Come scrivere un bel libro

Paperino Paolino^{∗} Paperopoli Gastone Paperone^{†} Fortulandia

Bologna, April 9, 2003

Abstract

Questo è un sommario

1 Prima sezione

1.1 Prima sottosezione

1.1.1 Prima sottosottosezione

2 Seconda sezione

2.1 con sottosezione

2.1.1 con sottosottosezione

∗Papero sfortunato

†Papero fortunato